OpenStudy (riley):

I'm doing something wrong with these derivative quotients and it's starting to really frustrate me. Calculate dy/dx. You need not expand your answer. 1) y = (7x^2-9x+11)/2x+4 The answer I keep coming up with which is wrong: dy/dx = (14x^2-38x+22)/(2x+4)^2 2) y = (x+2)(x+1)/2x-1 My answer, which is also wrong . . . dy/dx = (3x^2+x-5)/(2x-1)^2 Need some help please. =/

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{(7x^2-9x+11)'(2x+4)-(2x+4)'(7x^2-9x+11)}{(2x+4)^2}\]

myininaya (myininaya):

this is the first step are you ok with this?

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{(14x-9)(2x+4)-(2)(7x^2-9x+11)}{(2x+4)^2}\]

OpenStudy (riley):

Yes, that's how I'm plugging it in. Errr well I wrote it down like this: (14x)(2x+4) - (7x^2-9x+11)(2) and all of that over (2x+4)^2. Which I know the little ' sign means derivative of; am I doing that part wrong then?

OpenStudy (riley):

Oh oh oh the 9. >_< I completely dropped it.

OpenStudy (riley):

Lemme rewrite it all out again, two secs. That 9 was killing me . . .

myininaya (myininaya):

\[\frac{dy}{dx}=\frac{28x^2+56x-18x-36-14x^2+18x-22}{(2x+4)^2}\]

myininaya (myininaya):

ok

OpenStudy (riley):

Okay I'm not sure if this is right either . . . I re-worked it all over again and came up with: (14x^2-20x-58)/(2x+4)^2 O_O

OpenStudy (riley):

No thats not the right answer either. . . . amg what am I doing wrong!! =(

myininaya (myininaya):

i got a different coefficient of x above (56x-18x+18x)=56x

OpenStudy (riley):

. . . . .

OpenStudy (riley):

I didn't carry over one of my - signs . . .

OpenStudy (riley):

So then that leaves me with 14x^2+56x-58 / (2x+4)^2, I think.

myininaya (myininaya):

thats what i get

OpenStudy (riley):

lol, thats the right answer. . . gonna go bang my head against a wall really quick brb . .

OpenStudy (riley):

I'll go back and re-check my math for the second one. If i'm plugging it all in the correct way, then it has to be the simple math part that I'm screwing up on. Thank you for the help myininaya, it's very much appreciated.

myininaya (myininaya):

np